Multi-Variable Calculus Assignment

due Friday, February 9, 2001

Exercises

1. Find the locus of points equidistant from each of the three points (1, 0, 0), (0, 1, 0), and (0, 0, 1).

2. Find the projection of the vector [3, −6, 2] on the vector [1, 2, −2].

3. Find the distance of the point (4, −8, 3) from the line given parametrically by t ⟼ (1, −2, 1) + t[1, 2, −2] .

4. Find the distance of the point (4, −8, 3) from the plane given by the equation −x + 2y − 2z = 7 .

[Processed from GELLMU to L^{A}TEX on Tue 6 Feb 2001 at 15:31:30 EST]